Monday, 24 August

Plenary Lectures

9:00 – 10:00  Auditorium Maximum H 105

Karl Sigmund, University of Vienna, Vienna, Austria The Population Dynamics of Conflict and Cooperation (I: p.21)

Chair: Frank Hoppensteadt, Arizona State University, Tempe, USA

10:15 – 11:15  Auditorium Maximum H 105

Michel Talagrand, CNRS, Université Paris 6, Paris, France Huge Random Structures and Mean Field Models for Spin Glasses (I: p.23)

Chair: Lior Tzafriri, The Hebrew University, Jerusalem, Israel

11:45 – 12:45  Auditorium Maximum H 105

Cumrun Vafa, Harvard University, Cambridge, USA and Tehran, Iran Geometric Physics (I: p.25)

Chair: Maxim Kontsevich, IHES, Bures-sur-Yvette, France

ICM’98 – Program  1